# Time-Optimized Safe Navigation in Unstructured Environments through Learning Based Depth Completion

### v1

@@ 4.1 System Setup
For our experiments, we consider a $1.3\leavevmode\nobreak\ $\mathrm{k}\mathrm{g}$$ drone with a thrust to weight ratio of $4-1$. The onboard computer is an NVIDIA Orin NX communicating with a Pixracer Pro flight controller running PX4 [40]. All computation is performed onboard the system, and the station only handles visualization and simple initialization commands. This software stack is integrated with ROS2. Our sensing module is an Intel D455 RGB-D Camera. For localization we run either visual-inertial odometry algorithm OpenVINS [41] fused with IMU through an Unscented Kalman Filter (UKF) for indoor flight or GPS module for outdoor flight. Our monocular depth estimation network is DepthAnythingv2[16]. This model is optimized using TensorRT to ensure high speed inference at $80\leavevmode\nobreak\ $\mathrm{H}\mathrm{z}$$ inference time onboard. To solve our trajectory optimization in eq. (5), we use Interior Point Optimizer (IPOPT) [42] with MA57 [43] as the linear solver. We set a maximum velocity of $2\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}$$ per axis and similarly a max acceleration of $4\leavevmode\nobreak\ $\mathrm{m}\mathrm{/}\mathrm{s}^{2}$$ outdoors. This framework uses a SO(3) controller [44] to covert the optimized trajectory to rotation attitude commands for our quadrotor to fly through. During our exploration experiments, we use the Next Best View[39] algorithm to stress our navigation pipeline. Our Next Best View algorithm selects the goals with sample points sampled evenly across the entire map known and unknown regions. We conduct our tests in two areas an indoor flying space with a space of $8\leavevmode\nobreak\ $\mathrm{m}$\times 10\leavevmode\nobreak\ $\mathrm{m}$$, or a larger outdoor flying space $60\leavevmode\nobreak\ $\mathrm{m}$\times 40\leavevmode\nobreak\ $\mathrm{m}$$.
@@ 4.2 Outdoor Exploration
Figure 10: Trajectory Timing Comparison. Our Method’s Path planning time is under 1 $\mathrm{m}\mathrm{s}$ in all cases therefore is not visible in this plot. Similarly MINCO’s path planning in parking is around 1 $\mathrm{m}\mathrm{s}$ so not visible. SFC refers to the time for safe flight corridor generation. For the map names, we refer to Fig. 9.
Figure 11: Safe flight corridor generation. The height of these layers were manually set to prevent flying over obstacles.
In this set of experiments, we give the quadrotor exploration goal limits to reach $40\%-80\%$ of the declared volume seen in the flight path in Fig. 6. From no map, the quadrotor autonomously explores a region of $60\leavevmode\nobreak\ $\mathrm{m}$\times 40\leavevmode\nobreak\ $\mathrm{m}$$. Overall, we observe that as the planner is tasked with exploring a greater portion of the map, it naturally expands its search path, resulting in a more comprehensive coverage of the environment. Additionally, the vehicle is able to safely navigate in a collision-free manner through the environment in multiple trials showing the repeatability of of the proposed solution. When we combine all the trial flight times, they have a cumulative $20\leavevmode\nobreak\ $\mathrm{m}\mathrm{i}\mathrm{n}$$ of safe collision-free navigation in this environment. It took around $400\leavevmode\nobreak\ $\mathrm{s}$$ for the quadrotor to explore $80\%$ of the environment and about $300\leavevmode\nobreak\ $\mathrm{s}$$ to explore $40\%$. We can analyze the map generated when achieving $80\%$ exploration of the overall space based on its quality in Fig. 7. The map contains all buildings and maintains a similar geometric shape. Additionally, while we observe the presence of some speckle noise in the air, this amount remains minimal. Our tracking error and position following are close to the quadrotor desired value maintaining a tracking error of around $0.2\leavevmode\nobreak\ $\mathrm{m}$$. Although some positional error spikes are observed, they are primarily attributed to environmental factors, as the weather on the testing day included an average wind speed of $15\leavevmode\nobreak\ $\mathrm{m}\mathrm{p}\mathrm{h}$$ with strong $50\leavevmode\nobreak\ $\mathrm{m}\mathrm{p}\mathrm{h}$$ gusts.
@@ 4.3 Depth Generation Comparison
Additionally, we demonstrate that our learning-based depth performs far better than the commercial off-the-shelf depth generated by the D455 Stereo Cameras. First, the noise is reduced in the monocular version compared to the stereo. Next, we have an improved effective range in our depth completion. Third, we have a more complete image where structure is preserved. These three properties can be seen based on our experiments in Fig. 8 where we compare the point cloud from our stereo depth camera, Intel D455 which uses a stereo matching algorithm, and our depth completion method. Looking closer at Fig. 8, we notice that the noise is reduced in the left image with the stairs. Depth completion maintains a substantially better structure on the guard rails, stairs, and door. In contrast, the point cloud generated by the D455 depth has a wavelike noise that corrupts the depth’s structure. Additionally, because the walls are self-similar, we see holes in the depth estimates, especially next to the door and bottom window areas that are erroneously considered free space. None of these imperfections exist in the depth completion method proposed in this work. Second, we can see that the range of our depth completion method is improved in the middle image. The initial blue stereo depth only contains the beginning of the L building. On the other hand, because there is enough information on the scale from the initial depth, the depth completion can successfully interpolate the remainder of the building, creating the full structure and $90^{\circ}$ turn. Third, we compare the depth generated in the park. The depth completion point cloud is more complete preserving a greater amount of details. This includes the leaves of the trees, and much more of the park’s buildings are maintained. These features are not contained in the stereo baseline. Additionally, we show in our multimedia supplementary material a full video of the depth completion compared to the stereo depth over a $6\leavevmode\nobreak\ $\mathrm{m}\mathrm{i}\mathrm{n}$$ continuous walk. We notice that not only is the structure of our depth completion good, but also that the scale of the stereo depth image is consistent with depth completion in their common range. In our video, we additionally show that the generated depth is generalizable to both indoor and outdoor scenarios as we take the camera out of the basement. We also timed the average speed of our method as $1.5\leavevmode\nobreak\ $\mathrm{m}\mathrm{s}$$ on the Orin NX showing that not only is the quality improved, but our processing can be done in real-time onboard. Overall, these experiments demonstrate the superiority of our method compared to a standard, product-based stereo vision baseline.
@@ 4.4 Trajectory Planning Comparison
We compare our computation time and trajectory path with MINCO [11] which is a well-known global planner used in many similar works for obstacle avoidance works [2, 15] that creates obstacle-free and time-optimized global trajectories. Additionally MINCO [47] has shown to be more computationally efficient and optimal compared to multiple baselines in literature making it an effective benchmark. We compare our methods over $5$ separate maps some taken from popular works [11, 7, 46]. All computation timing and qualitative analysis are performed on an Orin NX to ensure a fair comparison. We record our findings in Fig. 9. Our pipeline takes less time to compute global obstacle free paths especially for complicated maps. This is especially true for the forest and hospital cases where we only take a third of the time of MINCO. Additionally, we can analyze the quality of our paths. While we often take a slightly longer path, our method maintains a similar path length quality to MINCO and contains safety constraints.
We analyze the breakdown of our trajectory planning times in Fig. 10. This includes a path planning, safe flight corridor generation, and finally a trajectory optimization. In our work, we measure the time needed to decompose the map as our safe flight corridor generation, and path planning is the amount of time needed to select among this graph the correct corridors. The results of safe flight corridor generation are shown in Fig. 11. We see the obstacles in gray and the layers of the cuboids in red, green, and blue. The free space is broken down to represent a tight fit around the obstacles. Our work initially assumes that unexplored space is a free space until an obstacle is observed. In MINCO’s, path planning is done using RRT implemented by [48]. Unlike our method, MINCO’s implementation generates SFC along the RRT path alone. In the timing breakdown, our work’s path planning takes under $1\leavevmode\nobreak\ $\mathrm{m}\mathrm{s}$$ making it invisible in Fig. 10 while MINCO path planning takes most of its run time. Additionally, if we compare only the last stage, we notice that our trajectory optimization has a similar completion time. However, we guarantee safety constraints as opposed to MINCO that treats safety as an objective in an unconstrained optimization. Part of the superior speed is based on our trajectory optimization’s objective to minimize the number of optimization variables. MINCO’s path planning objective is based on minimum distance instead. Overall, we show that our method produces faster global trajectories, maintains safety constraints, and is of similar quality to MINCO.

### v2

@@ 4.1 System Setup
For our experiments, we consider a $1.3\penalty 10000\ $\mathrm{k}\mathrm{g}$\$ drone with a thrust to weight ratio of 414-1. The onboard computer is an NVIDIA Orin NX communicating with a Pixracer Pro flight controller running PX4 [40]. All computation is performed onboard the system, and the station only handles visualization and simple initialization commands. This software stack is integrated with ROS2. Our sensing module is an Intel D455 RGB-D Camera. For localization we run either visual-inertial odometry algorithm OpenVINS [41] fused with IMU through an Unscented Kalman Filter (UKF) for indoor flight or GPS module for outdoor flight. Our monocular depth estimation network is DepthAnythingv2[15]. This model is optimized using TensorRT to ensure high speed inference at $80\penalty 10000\ $\mathrm{H}\mathrm{z}$\$ inference time onboard. To solve our trajectory optimization in eq. (5), we use Interior Point Optimizer (IPOPT) [42] with MA57 [43] as the linear solver. We set a maximum velocity of $2\penalty 10000\ $\mathrm{m}\mathrm{/}\mathrm{s}$\$ per axis and similarly a max acceleration of $4\penalty 10000\ $\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$ outdoors. This framework uses a SO(3) controller [44] to covert the optimized trajectory to rotation attitude commands for our quadrotor to fly through. During our exploration experiments, we use the Next Best View[39] algorithm to stress our navigation pipeline. Our Next Best View algorithm selects the goals with sample points sampled evenly across the entire map known and unknown regions. We conduct our tests in two areas an indoor flying space with a space of $8\penalty 10000\ $\mathrm{m}$\times 10\penalty 10000\ $\mathrm{m}$\$, or a larger outdoor flying space $60\penalty 10000\ $\mathrm{m}$\times 40\penalty 10000\ $\mathrm{m}$\$.
@@ 4.2 Outdoor Exploration
Figure 11: Safe flight corridor generation. The height of these layers were manually set to prevent flying over obstacles. The colors represent the various layers of the cuboid decomposition. Red color is the top layer. Green is the middle layer. Blue is the bottom layer
In this set of experiments, we give the quadrotor exploration goal limits to reach $40\%-80\%$ of the declared volume seen in the flight path in Fig. 7. From no map, the quadrotor autonomously explores a region of $60\penalty 10000\ $\mathrm{m}$\times 40\penalty 10000\ $\mathrm{m}$\$. Overall, we observe that as the planner is tasked with exploring a greater portion of the map, it naturally expands its search path, resulting in a more comprehensive coverage of the environment. Additionally, the vehicle is able to safely navigate in a collision-free manner through the environment in multiple trials showing the repeatability of of the proposed solution. When we combine all the trial flight times, they have a cumulative $20\penalty 10000\ $\mathrm{m}\mathrm{i}\mathrm{n}$\$ of safe collision-free navigation in this environment. It took around $400\penalty 10000\ $\mathrm{s}$\$ for the quadrotor to explore $80\%$ of the environment and about $300\penalty 10000\ $\mathrm{s}$\$ to explore $40\%$. We can analyze the map generated when achieving $80\%$ exploration of the overall space based on its quality in Fig. 8. The map contains all buildings and maintains a similar geometric shape. Additionally, while we observe the presence of some speckle noise in the air, this amount remains minimal. Our tracking error and position following are close to the quadrotor desired value maintaining a tracking error of around $0.2\penalty 10000\ $\mathrm{m}$\$. Although some positional error spikes are observed, they are primarily attributed to environmental factors, as the weather on the testing day included an average wind speed of $15\penalty 10000\ $\mathrm{m}\mathrm{p}\mathrm{h}$\$ with strong $50\penalty 10000\ $\mathrm{m}\mathrm{p}\mathrm{h}$\$ gusts.
@@ 4.3 Depth Generation Comparison
Method Name Method Type Residential KITTI (RMSE) Town KITTI (RMSE) Scene Flow Kitti (RMSE)
Stereo Block Matching Stereo Classical Vision $0.062$ $0.034$ $0.050$
Metric DepthAnything Monocular NN $0.57$ $0.21$ $0.12$
Our Method Linear Hybrid $0.056$ $0.030$ $\mathbf{0.012}$
Our Method Quadratic Hybrid $\mathbf{0.053}$ $\mathbf{0.029}$ $0.013$
Our Method Cubic Hybrid $0.054$ $0.030$ $0.013$
Table 3: Depth Completion Comparison on Kitti Dataset[47]. We use the inverted depth RMSE for our error metric. All error values are in $\mathrm{p}\mathrm{x}$ using the disparity RMSE with normalized focal length and baseline. Stereo Block Matching is a standard disparity method used in many off the shelf stereo cameras. Depth Anything Metric is the pure learning monocular depth method using image transformer which adds an additional linear head to rescale the image to a depth image. Hybrid refers to our combination of image transformer scaled with classical computer vision. Our Method linear, square and cube represents polynomial fitting for order 11, 22, and 33 polynomials respectively. For our images, we use Our Method *. The bold number is the minimum RMSE in the column.
Figure 12: The top image is the color input from the left camera of the stereo pair. The bottom four images show disparity maps produced by different methods. LiDAR provides the ground truth, obtained by projecting LiDAR depth measurements into the left camera frame. Depth Completion is our hybrid approach; the example shown here uses a second-order polynomial, though results are similar across different orders. Metric DepthAnythingv2[15] is a monocular learning-based method, with its output converted to disparity. Stereo Disparity is computed using block matching between the left and right stereo images.
To further validate our system, we evaluate our depth completion method. Compared to the commercial off-the-shelf Intel D455 stereo camera, our approach demonstrates three key advantages: reduced noise, an extended effective range, and improved structural consistency. These improvements are evident in Fig. 9, which compares point clouds generated by the D455 stereo camera, a classical computer vision block matching algorithm, against our depth completion results. In the double story building example, the stereo reconstruction exhibits strong wavelike noise that degrades the geometry of the guard rails, stairs, and door. In contrast, our depth completion significantly reduces noise and preserves fine structural details. Additionally the stereo system struggles with self-similar surfaces, the D455 depth maps frequently contain holes, such as around the door and window regions. These artifacts are largely eliminated in our method. Moreover, the effective range of our approach is improved. For instance, in the middle image, the stereo baseline only partially reconstructs the L-shaped building. Leveraging the scale from stereo depth, our completion algorithm interpolates the missing structure, correctly reproducing the building’s full geometry and 90° corner. Finally, when tested in the park environment, our method generates denser and more complete reconstructions, retaining fine details such as foliage and preserving much more of the surrounding buildings than the stereo baseline. Additional qualitative results are shown in our video. Depth completion is generalizable to both indoor and outdoor scenarios as we take the camera out of the basement. We also timed the average speed of our method as $1.5\penalty 10000\ $\mathrm{m}\mathrm{s}$\$ on the Orin NX showing that not only is the quality improved, but our processing can be done in real-time onboard.
We further quantify the performance of our depth generation method using the KITTI dataset[47]. To emphasize accuracy in the near field—critical for collision avoidance—we evaluate disparity root-mean-square error (RMSE) rather than depth RMSE, since disparity naturally weights errors at closer ranges more heavily. The error is computed as:
$\frac{1}{A_{m}}\sum_{j=0}^{W}\sum_{i=0}^{H}\left(\mathbf{M}[i,j]\tilde{\mathbf{d}}[i,j]-(\mathbf{M}[i,j]\mathbf{d}[i,j]\right)^{2}.$ (10)
$\tilde{\mathbf{d}}$ is the ground truth disparity gathered from the LiDAR depth in the KITTI dataset[47]. $\mathbf{d}$ is the tested method’s disparity. $\mathbf{d}$ is generated from either a single or stereo color image gathered from the cameras. $\mathbf{M}$ is a mask image representing the valid disparities in both the ground truth and generated disparity image. $A_{m}$ is the number of valid pixels in in the mask image and the estimated depth image, $\mathbf{d}$. LiDAR data is only used as a ground truth reference. Table. 3 reports the root mean square error (RMSE) across multiple KITTI scenes and various polynomial orders. Among all methods, our hybrid depth completion consistently achieves the lowest error. In particular, the second-order (quadratic) polynomial fitting yields the most accurate results overall, although the first- and third-order variants perform comparably. By contrast, the pure monocular learning method (Metric DepthAnything) performs the worst due to the sim-to-real gap. While the strucutre is similar to our depth completion method, the scale is on the whole poor as see in Fig. 12. Standard stereo block matching achieves reasonable performance but is limited by structural noise and missing regions, as reflected in both the qualitative and quantitative results. These missing regions makes it unsuitable for use in a mapping context.
@@ 4.4 Trajectory Planning Comparison
[…]
Town $12000$ $0.020$ <0.001<0.001 $0.005$ $0.043$
Hospital $16000$ $0.101$ <0.001<0.001 $0.600$ $0.043$
Parking $500$ $0.006$ <0.001<0.001 <0.001<0.001 $0.014$
Forest $1000$ $0.015$ <0.001<0.001 $0.200$ $0.025$
Perlin Noise $12500$ $0.075$ <0.001<0.001 $0.076$ $0.040$
Table 4: Path Planning Timing: Cuboid Decomp is our method which is the time it takes to covert the occupancy map to a graph of convex sets. We use A* on Convex Sets to select the optimal set of safe flight corridors. RRT is rapidly exploring random tree implemented using the OMPL toolbox[48] and used in MINCO as an alternate path planning. Convex Polytope is the Safe Flight Corridor generation method over RRT used in MINCO.
Map Size $30\%$ Decomp. Time (s)($\mathrm${s}\$) $30\%$ A* Time (s)($\mathrm${s}\$) $50\%$ Decomp. Time (s)($\mathrm${s}\$) $50\%$ A* Time (s)($\mathrm${s}\$) $65\%$ Decomp. Time (s)($\mathrm${s}\$) $65\%$ A* Time (s)($\mathrm${s}\$)
$25$\mathrm{m}$\times 25$\mathrm{m}$\times 5$\mathrm{m}$\$ $0.075$ <0.001<0.001 $0.14$ <0.001<0.001 $0.18$ <0.001<0.001
$50$\mathrm{m}$\times 50$\mathrm{m}$\times 5$\mathrm{m}$\$ $0.542$ $0.003$ $0.689$ $0.003$ $0.851$ $0.004$
$75$\mathrm{m}$\times 75$\mathrm{m}$\times 5$\mathrm{m}$\$ $1.75$ $0.014$ $2.01$ $0.017$ $2.24$ $0.018$
Table 5: Cuboid Decomposition timing on each map is randomly generated using Perlin Noise for different infill levels and different volumes. The A* Timing is the average timing it takes to navigate one corner of the map to another.
We benchmark our trajectory planning framework against MINCO [29], a widely used global planner known for producing time-optimal, collision-free trajectories with strong computational efficiency. All evaluations are performed onboard the NVIDIA Orin NX to ensure fair comparison. We demonstrate our results in Fig. 10 compares paths generated by both methods across five representative maps drawn from prior works [29, 20, 46]. Our trajectories are generally comparable in length to those produced by MINCO, though occasionally slightly longer. However, our method explicitly enforces safety constraints, whereas MINCO does not guarantee safety and can return potentially unsafe trajectories. Fig. 11 illustrates the safe flight corridors generated by our cuboid decomposition method. By tightly fitting free space around obstacles, we ensure robust feasibility even in cluttered environments and improved computation time.
We show this computation time in Table 4 which presents a timing breakdown for different planning components. Our pipeline consists of two main steps: (i) cuboid decomposition of the occupancy map into a graph of convex sets, and (ii) A* search over this graph to select the optimal sequence of safe flight corridors. Both steps are highly efficient: decomposition can be solved in parallel before the planning occurs, and A* search completes in under 1 ms. In contrast, MINCO relies on RRT-based sampling for path discovery, which dominates its runtime. Consequently, across all test maps, our framework computes global trajectories substantially faster than MINCO as seen in Fig. 10. Fig. 10 contains the full trajectory pipeline computation time.
Finally, we evaluate the systems scalability. Table 5 evaluates runtime scalability under increasing map sizes generated by Perlin noise. As expected, cuboid decomposition time scales approximately linearly with environment volume. While this may limit update frequency in very large-scale maps, the subsequent A* search remains negligible compared to decomposition. In summary, our trajectory planning approach consistently achieves lower computation times than MINCO while preserving comparable path quality and offering guaranteed safety constraints. These properties make it well-suited for real-time autonomous flight in cluttered and dynamically evolving environments.
@@ 5 Conclusion
APPENDIX
This section details the definition for the Bernstein Polynomials and explain the derivations for eq. 4 6 7. All properties are derived from the basis function, eq. 11 . In this section we consider a single spline of order $n$ and time duration $T$.
$\begin{split}b_{v,n}(t)&=\binom{n}{v}\left(\frac{t}{T}\right)^{v}\left(1-\frac{t}{T}\right)^{n-v}\\ B_{n}(t)&=\sum_{v=0}^{n}c_{v,n}b_{v,n}(t)\penalty 10000\ t\in[0,T]\end{split}$ (11)
where $v$ represents the basis or coefficient number $t$ represents time and $\binom{n}{v}=\frac{n!}{(n-v)!v!}$. Because this section needs to go over the generalized form of Bernstein Polynomial for any order, we change the definition of our coefficient subscript from above $c_{v,n}$. The following represents the $v$ coefficient of a $n$ order Bernstein polynomial. Using the definition of the basis function the following properties, start and end point special definitions, derivative definition, and the convex hull property will be derived.
5.1 Bernstein Polynomial Derivative
In this section, we will prove that the time derivative represented by the symbol for our Bernstein polynomial of order $n$ can be represented by a lower order n1n-1 Bernstein polynomial with the following coefficients
$B^{\prime}_{n}(t)=\frac{n}{T}\sum_{v=0}^{n-1}b_{v,n-1}(t)\left(c_{v+1,n}-c_{v,n}\right).$ (12)
We first take the derivatives of our basis function and get $3$ cases, $v=0,n$ and $v\in[1,n-1]$
$\begin{split}\text{$b^{\prime}_{0,n}\left(t\right)$}&=-\frac{n}{T}\left(b_{0,n-1}(t)\right),\\ \text{$b^{\prime}_{v\in[1,n-1],n}\left(t\right)$}&=\frac{n}{T}\left(b_{v-1,n-1}(t)-b_{v,n-1}(t)\right),\\ \text{$b^{\prime}_{n,n}\left(t\right)$}&=\frac{n}{T}\left(b_{n-1,n-1}(t)\right).\end{split}$ (13)
This result comes from the rewriting the basis function as
[…]
We can then write the time derivative as the follow values.
$b^{\prime}_{v,n}(t)=\binom{n}{v}f^{\prime}(t)g(t)+\binom{n}{v}f(t)g^{\prime}(t)$ (15)
We can derive the first term as
$\binom{n}{v}f^{\prime}(t)g(t)=\begin{cases}\text{$0$}&\quad\text{if $v=0$}\\ \text{$\frac{n}{T}b_{v-1,n-1}(t)$}&\quad\text{if $v>0$}\end{cases}.$ (16)
The second term forms
$\binom{n}{v}f(t)g^{\prime}(t)=\begin{cases}\text{$0$}&\quad\text{if $v=n$}\\ \text{$-\frac{n}{T}b_{v,n-1}(t)$}&\quad\text{if $v<n$}\end{cases}.$ (17)
When we combine these two terms, we get the $3$ special cases in eq. (13). By substituting the time derivative of the basis functions in our Bernstein polynomial we get the following summation
$\begin{split}B^{\prime}_{n}(t)=\frac{n}{T}(c_{n,n}b_{n-1,n-1}(t)\\ +\frac{n}{T}\sum_{v=1}^{n-1}c_{v,n}\left(b_{v-1,n-1}(t)-b_{v,n-1}(t)\right)\\ -\frac{n}{T}c_{0,n}b_{0,n-1}(t),\end{split}$ (18)
We can realize this equation is exactly the same as the eq. (12). We can represent this using matrix multiplication as follows
[…]
We can extrapolate this to a higher order time derivatives by repeatedly applying eq. (19). This results form the core of eq. (7).
5.2 Bernstein Endpoints
Let us recall, the basis function is constructed from products of $\frac{t}{T}$ and $1-\frac{t}{T}$. This means for the special cases $B_{n}(0)$ and $B_{n}(T)$. The basis function is a result of powers of $1$ and 0, This property enforces that the only active basis functions are the ones with an active $0^{0}$ power. This gives the following properties.
$B_{n}(0)=c_{0,n},\penalty 10000\ \penalty 10000\ B_{n}(T)=c_{n,n}.$ (20)
Using eq. (19), we substitute time values 0 and $T$ to get the following
$\begin{split}B^{\prime}(0)&=\frac{n}{T}(c_{1,n}-c_{0,n}),\\ B^{\prime}(T)&=\frac{n}{T}(c_{n,n}-c_{n-1,n}).\end{split}$ (21)
Finally, we use the same logic to acquire the relationship between the start and end acceleration, ′′, of the polynomial as
$\begin{split}B^{\prime\prime}(0)&=\frac{n(n-1)}{T^{2}}(c_{2,n}-2c_{1,n}+c_{0,n}),\\ B^{\prime\prime}(T)&=\frac{n(n-1)}{T^{2}}(c_{n,n}-2c_{n-1,n}+c_{n-2,n}).\end{split}$ (22)
We can simply stack the matrices together and get the following result for an order $5$ polynomial
$\begin{bmatrix}B(0)\\ B^{\prime}(0)\\ B^{\prime\prime}(0)\\ B(T)\\ B^{\prime}(T)\\ B^{\prime\prime}(T)\end{bmatrix}=\begin{bmatrix}1&0&0&0&0&0\\ -\frac{5}{T}&\frac{5}{T}&0&0&0&0\\ \frac{20}{T^{2}}&-\frac{40}{T^{2}}&\frac{20}{T^{2}}&0&0&0\\ 0&0&0&0&0&1\\ 0&0&0&0&-\frac{5}{T}&\frac{5}{T}\\ 0&0&0&\frac{20}{T^{2}}&-\frac{40}{T^{2}}&\frac{20}{T^{2}}\end{bmatrix}\begin{bmatrix}c_{0,5}\\ c_{1,5}\\ c_{2,5}\\ c_{3,5}\\ c_{4,5}\\ c_{5.5}\end{bmatrix}$. (23)
We acquire eq. (4) by inverting the above equation. This forms the relationship between our waypoints and the order $5$ Bernstein polynomial.
5.3 Bernstein Convex Hull Property
Our goal is to prove the following theory where $l$ and $u$ are constant lower and upper bounds
$\begin{bmatrix}l\\ \vdots\\ l\end{bmatrix}<\begin{bmatrix}c_{0,n}\\ \vdots\\ c_{n,n}\end{bmatrix}<\begin{bmatrix}u\\ \vdots\\ u\end{bmatrix}\Rightarrow l<B_{n}(t)<u,$ (24)
Proving this property allows us to represent a continuous time constraint using only $n$ discrete constraints. Additionally, since eq. 4 establishes the waypoint and coefficient relationship we can combine eq. (4) and eq. (24) to obtain eq. (2).
To prove this property is true, we perform the following steps First, we sum our vectors on each side.
$\begin{bmatrix}l\\ \vdots\\ l\end{bmatrix}\leq\begin{bmatrix}c_{0,n}\\ \vdots\\ c_{n,n}\end{bmatrix}\leq\begin{bmatrix}u\\ \vdots\\ u\end{bmatrix}\Rightarrow\sum_{v=0}^{n}l\leq\sum_{v=0}^{n}c_{v,n}\leq\sum_{v=0}^{n}u,$ (25)
We then multiple each term of the summation individually by the basis function $b_{v,n}$. Because each basis function is always >0>0 by its construction as we bind the polynomial between $t\in[0,T]$, this implies
$l\sum_{v=0}^{n}b_{v,n}(t)\leq\sum_{v=0}^{n}c_{v,n}b_{v,n}(t)\leq u\sum_{v=0}^{n}b_{v,n}(t).$ (26)
We would like invoke the binomial theorem from algebra [49] and the basis function, eq. 11, to derive that the sum of our basis function is equal to $1$.
$\begin{split}(x+y)^{n}&=\sum_{v=0}^{n}\binom{n}{v}x^{v}y^{n-v},\\ b_{v,n}(t)&=\binom{n}{v}\left(\frac{t}{T}\right)^{v}\left(1-\frac{t}{T}\right)^{n-v},\\ \sum_{v=0}^{n}b_{v,n}(t)&=\left(1-\frac{t}{T}+\frac{t}{T}\right)^{n}=1.\end{split}$ (27)
Using these values, we are able formulate the final part of our proof by replacing $\sum_{v=0}^{n}b_{v,n}=1$ in eq. (26)
$l\leq\sum_{v=0}^{n}c_{v,n}b_{v,n}(t)\leq u\Rightarrow l\leq B_{n}(t)\leq u.$ (28)
This means that if the coefficients are bound by a lower and upper value then the resulting function is also bound by the same coefficients. We can apply the same properties to bind the future derivatives. This forms the proof for eq. 6 and allows us to set continuous time constraints using a finite number of discrete constraints. We can also apply this to higher order derivative constraints by combining this property with the derivative property.
5.4 Distance Metrics for Convex Sets
We will motivate why distance metric between sets is difficult to define in Fig. 14. Formally, Fig. 14 is defined using the following set of $2-\text{D}$ convex corridors. Each corridor is in the form $x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]$
  • Corridor 1: $4,-3]\times[-2,2]$ Center (3.5,0)(-3.5,0)
  • Corridor 2: $1,1]\times[-2,2]$ Center $(0,0)$
  • Corridor 3: $5,5]\times[0,1]$ Center $(0,0.5)$
  • Corridor 4: $6,-4.5]\times[0,2]$ Center (5.25,1)(-5.25,1)
  • Target: (5.1,0.5)(-5.1,0.5).
We will demonstrate analyze following metrics below
Figure 14: The top map is a drawing of how the convex sets intersect. The bottom graphs are their abstracted from with edge weights. The green value represents the heuristic. The Star represents the target.
Shortest Distance between two sets. This metric is useful for a heuristic for long term planning, but we see the definition of two convex sets being connected this value is always zero. We see that calculating the shortest distance between a set and the target gives good guidance for each set. This makes it unsuitable for a search metric additionally there becomes no penalty for including an arbitrary number of convex sets.
Constant Edge Weight. This minimizes the number of splines or optimization variables as we naturally are minimizing the number of safe flight corridors. This does not minimize the minimum distance as we see the paths 1341-3-4 is equal to path 2342-3-4. We add the heuristic metric previously to encode some amount of optimality between two sets. The path 1341-3-4 becomes a shorter path. We found in our experimental results this is a good approximation for the maps we face.
Center of Cuboids. This metric does not have a useful physical meaning in distance. Corridors closer to the center are considered lower distance than corridors farther from the center. The distance from corridor 3 to 4 is further than the distance from the corridor 1 to 3. As a result, we see path 1341-3-4 has a much longer distance than path 2342-3-4. This is the opposite of reality. Adding the heuristic to the search still has the incorrect path 2342-3-4 be further.